# Unveiling the Role of ChatGPT in Software Development: Insights from Developer-ChatGPT Interactions on GitHub

### v1

@@ 6. Threats to Validity
Construct validity: Potential threats of this study lie in the mined ChatGPT links and their content that developers shared on GitHub. Not all shared ChatGPT links reflect meaningful interactions, and some may be noise, off-topic, or duplicates. For example, many shared ChatGPT links are not relevant to software development, or certain development-related ChatGPT links are duplicated from different sources (e.g., one link occurs in both Code and Issues). To mitigate these threats, we employed rigorous data cleaning techniques to eliminate noise data (e.g., irrelevant and duplicates) and ensure the uniqueness of mined data through a specific ID filtering process.
External validity: A relevant threat to the generalizability of our findings concerns the temporal stability of the observed findings. Our analysis focuses on developer-ChatGPT interactions on GitHub from May 2023 to June 17, 2024, a period that coincides with the release of ChatGPT’s sharing feature. As GenAI models and their applications continue to evolve, the patterns of usage and perceived effectiveness may also shift over time. Additionally, the findings may not fully generalize to other open-source platforms or proprietary development environments, which may involve different user demographics or AI tools. While our study captures a significant portion of ChatGPT usage among developers, it may not fully represent the broader spectrum of usage across all development contexts. Nevertheless, given GitHub’s prevalence as a leading platform for collaborative software development, we believe that our findings reasonably reflect broader trends within the developer community.
@@ 7. Conclusions and Future Work
Our findings reveal the characteristics of ChatGPT’s usage on GitHub, such as ChatGPT’s prompt turns distribution, and accompanying description. Moreover, the results uncover five primary purposes for which developers share developer-ChatGPT conversations, with Task Delegation being the dominant purpose, primarily to automate work. Furthermore, our results emphasize the widespread adoption of ChatGPT across nine development-related activities, with Code and Commits serving as the main data sources, and Software Development, Software Maintenance and Evolution as the dominant activities in which developers frequently utilize ChatGPT. In addition, we established a comprehensive mapping framework to analyze the specific SE tasks for which developers employ ChatGPT during their daily routines. Overall, our study demonstrates the pivotal role of GenAI tools represented by ChatGPT in supporting a wide range of SE activities across the SDLC.
@@ Acknowledgements.
- Hao et al. (2024) Huizi Hao, Kazi Amit Hasan, Hong Qin, Marcos Macedo, Yuan Tian, Steven HH Ding, and Ahmed E Hassan. 2024. An Empirical Study on Developers Shared Conversations with ChatGPT in GitHub Pull Requests and Issues. Empirical Software Engineering 29, 6 (2024), 150.
- Hou et al. (2023) Xinyi Hou, Yanjie Zhao, Yue Liu, Zhou Yang, Kailong Wang, Li Li, Xiapu Luo, David Lo, John Grundy, and Haoyu Wang. 2023. Large language models for software engineering: A systematic literature review. ACM Transactions on Software Engineering and Methodology 33, 8 (2023), 1–79.
@@
- Li et al. (2025) Ruiyin Li, Peng Liang, Yifei Wang, Yangxiao Cai, Weisong Sun, and Zengyang Li. 2025. Dataset for the Paper: Unveiling the Role of ChatGPT in Software Development: Insights from Developer-ChatGPT Interactions on GitHub.

### v2

@@ 6. Threats to Validity
Construct validity: Potential threats of this study lie in the mined ChatGPT links and their content that developers shared on GitHub. Not all shared ChatGPT links reflect meaningful interactions, and some may be noise, off-topic, or duplicates. For example, many shared ChatGPT links are not relevant to software development, or certain development-related ChatGPT links are duplicated from different sources (e.g., one link occurs in both Code and Issues). To mitigate these threats, we employed rigorous data cleaning techniques to eliminate noisy data (e.g., irrelevant and duplicates) and ensure the uniqueness of mined data through a specific ID filtering process.
External validity: A relevant threat to the generalizability of our findings concerns the temporal stability of the observed findings. Our analysis focuses on developer-ChatGPT interactions on GitHub from May 2023 to June 17, 2024, a period that coincides with the release of ChatGPT’s sharing feature. As GenAI models and their applications continue to evolve, the patterns of usage and perceived effectiveness may also shift over time. Additionally, the findings may not fully generalize to other open-source platforms or proprietary development environments, which may involve different user demographics or AI tools. While our study captures a significant portion of shared ChatGPT usage among developers on GitHub, it may not fully represent the broader spectrum of ChatGPT usage across all development contexts. Nevertheless, given GitHub’s prevalence as the leading platform for collaborative software development, we believe that our findings reasonably capture the prevailing trends within the developer community.
@@ 7. Conclusions and Future Work
Our findings reveal the characteristics of ChatGPT usage on GitHub, such as ChatGPT’s prompt turns distribution, and accompanying description. Moreover, the results uncover five primary purposes for which developers share developer-ChatGPT conversations, with Task Delegation being the dominant purpose, primarily to automate work. Furthermore, our results emphasize the widespread adoption of ChatGPT across nine development-related activities, with Code and Commits serving as the main data sources, and Software Implementation, Software Maintenance and Evolution as the dominant activities in which developers frequently utilize ChatGPT. In addition, we established a comprehensive mapping framework to analyze the specific SE tasks for which developers employ ChatGPT during their daily routines. Overall, our study demonstrates the pivotal role of GenAI tools represented by ChatGPT in supporting a wide range of SE activities across the SDLC.
@@ Acknowledgements.
- AlOmar et al. (2024) Eman Abdullah AlOmar, Anushkrishna Venkatakrishnan, Mohamed Wiem Mkaouer, Christian Newman, and Ali Ouni. 2024. How to refactor this code? an exploratory study on developer-ChatGPT refactoring conversations. In Proceedings of the 21st International Conference on Mining Software Repositories (MSR). 202–206.
@@
- Hao et al. (2024) Huizi Hao, Kazi Amit Hasan, Hong Qin, Marcos Macedo, Yuan Tian, Steven HH Ding, and Ahmed E Hassan. 2024. An Empirical Study on Developers’ Shared Conversations with ChatGPT in GitHub Pull Requests and Issues. Empirical Software Engineering 29, 6 (2024), 150.
- Hou et al. (2024) Xinyi Hou, Yanjie Zhao, Yue Liu, Zhou Yang, Kailong Wang, Li Li, Xiapu Luo, David Lo, John Grundy, and Haoyu Wang. 2024. Large language models for software engineering: A systematic literature review. ACM Transactions on Software Engineering and Methodology 33, 8 (2024), 1–79.
@@
- Li et al. (2025) Ruiyin Li, Peng Liang, Yifei Wang, Yangxiao Cai, Weisong Sun, and Zengyang Li. 2025. Dataset for the Paper: Unveiling the Role of ChatGPT in Software Development: Insights from Developer-ChatGPT Interactions on GitHub. https://github.com/RuiyinL/DevChat.
@@
- Luo et al. (2022) Xianchang Luo, Yinxing Xue, Zhenchang Xing, and Jiamou Sun. 2022. Prcbert: Prompt learning for requirement classification using bert-based pretrained language models. In Proceedings of the 37th IEEE/ACM International Conference on Automated Software Engineering (ICSE). 1–13.
@@
- Mining Challenge Track of MSR 2024 (2024) Mining Challenge Track of MSR 2024. 2024. Retrieved September 05, 2025 from https://2024.msrconf.org/track/msr-2024-mining-challenge.
- Mohamed et al. (2024) Suad Mohamed, Abdullah Parvin, and Esteban Parra. 2024. Chatting with AI: Deciphering developer conversations with ChatGPT. In Proceedings of the 21st International Conference on Mining Software Repositories (MSR). 187–191.
@@
- Rabbi et al. (2024) Md Fazle Rabbi, Arifa Islam Champa, Minhaz F Zibran, and Md Rakibul Islam. 2024. AI writes, we analyze: The ChatGPT Python code saga. In Proceedings of the 21st International Conference on Mining Software Repositories (MSR). 177–181.
@@
- Siddiq et al. (2024) Mohammed Latif Siddiq, Lindsay Roney, Jiahao Zhang, and Joanna Cecilia Da Silva Santos. 2024. Quality assessment of ChatGPT generated code and their use by developers. In Proceedings of the 21st International Conference on Mining Software Repositories (MSR). 152–156.
@@
- Wu et al. (2024) Liangxuan Wu, Yanjie Zhao, Xinyi Hou, Tianming Liu, and Haoyu Wang. 2024. ChatGPT chats decoded: Uncovering prompt patterns for superior solutions in software development lifecycle. In Proceedings of the 21st International Conference on Mining Software Repositories (MSR). 142–146.